# Metis-SPECS: Decoupling Learning via Self-distilled Preference-based Cold Start for VLMs

### v1

@@ 1 Introduction
Recently, inspired by the success of DeepSeek-R1 (Guo et al., 2025), in effectively enhancing the reasoning capabilities of large models through reinforcement learning (RL) with verifiable reward (Lambert et al., 2024; Guo et al., 2025), a growing body of work has begun to apply RL directly to vision language models (VLMs). This has led to a wave of exciting “MLLM-r1” research (Meng et al., 2025; Shen et al., 2025; Peng et al., 2025; Zhou et al., 2025; Zhang et al., 2025a; Wang et al., 2025b; a; Zheng et al., 2025; Ma et al., 2025), which leverage similar principles to advance multimodal reasoning.
@@
Previous research has indicated that prior to RL, employing a pre-training or warm-up phase (which is termed “cold start”), can significantly improve the readability, stability, and even the final performance of RL training (Guo et al., 2025). Currently, the most commonly used cold start strategy is supervised fine-tuning (SFT) , where the model is first fine-tuned on a set of high-quality Chain-of-Thought reasoning data to provide a better initial policy for the subsequent RL phase (Wei et al., 2025; Yang et al., 2025; Huang et al., 2025; Deng et al., 2025b). This strategy enables the model to be trained on complex reasoning data during the cold start phase, thereby acquiring reasoning ability.
@@
The common understanding behind SFT-based cold start is that reasoning abilities, reasoning format and other learning objectives can be jointly learned during the cold start phase. However, such an SFT-based joint learning paradigm may largely affect the model’s generalization capability (Wu et al., 2025; Chu et al., 2025), and consequently degrade subsequent RL (Chen et al., 2025a). This raises an important research issue of quantifying and improving the model’s generalization capability during cold start and working in concert with subsequent RL.
@@
To address the above limitations, we consider an alternative learning paradigm, which separates the learning process into hierarchical stages based on the idea that cold start phase focused more on shallow, surface-form learning to avoid prematurely getting stuck in in-distribution problem solving, while subsequent RL focuses on the deep-level learning of a solution to boost the overall performance (Bengio et al., 2009). Thus, the intuition of our adopting decoupling learning for multimodal reasoning is that the selection of pre-training methods in cold start needs to better support the subsequent RL, both in terms of generalization and by having separate objectives to facilitate better final results.
@@
Another important issue is the generation of cold start data. Previously, the prohibitive cost of human annotation has motivated a growing body of research to explore the use of synthetic data. This often involves using a more capable large model as a “teacher” to distill data for a smaller “student” model. (Zhang et al., 2025b; Yao et al., 2024; Xu et al., 2024; Huang et al., 2025). However, when the capability gap between the teacher model and the student model is too large, it can lead to a decline in model performance (Zhang et al., 2023). Alternatively, the DeepSeek-R1-Zero paradigm (Guo et al., 2025) first directly applies reinforcement learning to the base model for obtaining R1-Zero and then generates cold start data by zero model itself. This paradigm has achieved very remarkable performance; yet it still has the limitation of reliance on the SFT cold start and the constraints between SFT and subsequent RL, thereby leaving room for further improvement.
@@
In this paper, to examine the suitable cold start training method, we propose the Generalization Factor (GF) coefficient in Section 2 to quantify the generalization capability of the model and conduct an empirical study to evaluate different training methods. We identify that Direct Preference Optimization (DPO) (Rafailov et al., 2023) based on preference data is a cold start approach that enables the model to have better performance. On this basis, we present the Self-distilled Preference-based Cold-Start (SPECS) framework in Section 3. By decoupling the learning objectives during DPO to focus on output format, we create a pre-aligned model that serves as a superior starting point for the final RL fine-tuning. Our experiments show that this method leads to more stable, efficient training, and a higher performance ceiling compared to the advanced and strong baseline.
@@ 2.1 Evaluating Degree of Generalization
Setup. We define an evaluation function $\psi(f,P)\in\mathbb{R}$ that measures the performance of a model $f$ on a data distribution $P.$ A higher value of $\psi$ indicates better performance.
We consider two types of training data, denoted by $\tau\in\{pref,sup\}$, corresponding to preference data and supervised data. For a given data type $\tau$ and a fixed sample size $n$, a model $f_{\tau,n}$ is obtained by applying an algorithm $A_{\tau}$ to a training set $S_{\tau,n}$ which consists of $n$ samples drawn from a training distribution $P_{train}$.
$f_{\tau,n}=\mathcal{A}_{\tau}(S_{\tau,n}),\quad\mathrm{where}\quad S_{\tau,n}\sim P_{train}^{n}$
Generalization Factor. To assess the model’s performance, we measure its effectiveness on both in-distribution (ID) and out-of-distribution (OOD) data.
  • ID Performance: The ID performance, $\Psi_{\mathrm{ID}}^{\tau}(n)$, is evaluated on a hold-out set from the same distribution $P_{train}$
    $\Psi_{\mathrm{ID}}^{\tau}(n)=\psi(f_{\tau,n},P_{train})$
  • OOD Performance: The OOD performance, $\Psi_{\mathrm{OOD}}^{\tau}(n)$, is the weighted average performance across a set of $m$ distinct OOD distributions, $Q=\{Q_{1},\ldots,Q_{m}\}$, with weights defined by a distribution $\alpha$.
    $\Psi_{\mathrm{OOD}}^{\tau}(n)=\mathbb{E}_{Q\sim\alpha}[\psi(f_{\tau,n},Q)]$
We establish a baseline model, $f_{\tau,0}$, which serves as a reference point. This is typically a pre- trained model that has not been fine-tuned on data of type $\tau$. The performance gains over this baseline are calculated as:
$G_{\mathrm{ID}}^{\tau}(n)=\Psi_{\mathrm{ID}}^{\tau}(n)-\Psi_{\mathrm{ID}}^{\tau}(0)$
$G_{\mathrm{OOD}}^{\tau}(n)=\Psi_{\mathrm{OOD}}^{\tau}(n)-\Psi_{\mathrm{OOD}}^{\tau}(0)$
With a fixed number of samples $n$, we formalize the GF for data type $\tau$, denoted as $\Gamma_{\tau}(n)$, weighted harmonic mean of OOD performance gain and ID performance gain.
$\Gamma_{\tau}(n)=\frac{(1+\beta^{2})G_{\mathrm{OOD}}^{\tau}(n)G_{\mathrm{OOD}}^{\tau}(n)}{\beta^{2}G_{\mathrm{ID}}^{\tau}(n)+G_{\mathrm{ID}}^{\tau}(n)}$
@@ 2.2 Experimental findings
To preliminarily examine how preference data and supervised data affect model generalization, we construct a preference dataset $\mathcal{D}_{\mathrm{pref}}=\{(x_{i},y_{i}^{+},y_{i}^{-})\}_{i=1}^{N}$ , where $y_{i}^{+}$ is the chosen response and $y_{i}^{-}$ is the rejected response, and a supervised dataset $\mathcal{D}_{\mathrm{SFT}}=\{(x_{i},y_{i})\}_{i=1}^{N}$ with $y_{i}=y_{i}^{+}$ around reasoning tasks defined by a specific answer format. Under equal data budgets, we evaluate two settings: (i) an in-distribution setting in which the required reasoning format matches that used in training, and (ii) an out-of-distribution setting in which the required reasoning format differs  [^2] For more detailed description, see Appendix A.2. We compare DPO training, SFT training, and DPO training augmented with SFT loss (see Section 3.3). The resulting $G_{\mathrm{ID}}^{\tau}(n)$, $G_{\mathrm{OOD}}^{\tau}(n)$, and the $\Gamma_{\tau}(n)$ are reported in Figure 1.
@@ 3.2 Self-Distillation for Preference Data Generation
  • Chosen Response Filtration. For the chosen response ($y_{i}^{+}$), we use Gemini-2.5 flash (Comanici et al., 2025) as an evaluator. Assesses whether the reasoning path in the $\pi_{GRPO-zero}$ response aligns correctly with its final answer. Only responses in which the reasoning and the answer are consistent are retained, forming a high-quality pool of candidates.
@@ 3.3 DPO-based Pre-Alignment for Cold-Start
Methodology: To achieve this pre-alignment, we employ DPO (Rafailov et al., 2023), a powerful technique that directly optimizes the language model on preference data without the need for an explicit reward model. The standard DPO loss function is defined as:
@@
To augment this process, we incorporate an SFT loss computed on the “chosen” samples, which serves as a form of regularization. It ensures that while the model learns the directional preference signal from DPO, it does not drift far from the core distribution of high-quality text embodied by the chosen responses (Rao et al., 2025). The combined loss function is thus:
@@
where $\mathcal{L}_{SFT}$ is the conventional negative log-likelihood loss on the chosen responses, and $\lambda$ is a weighting coefficient to balance the two objectives. For a discussion for $\lambda$, see Appendix A.3.
@@ 3.4 Final GRPO Fine-tuning
For the final stage of fine-tuning, we employ the GRPO algorithm (Shao et al., 2024). This process is guided by a composite reward function that combines format and accuracy components to evaluate the model’s output, $o$, for a given question, $q$.
@@ 4.1 Experiment Settings
Dataset and Benchmark: The data utilized for training $\pi_{GRPO-zero}$ model in Stage 1 and for the final GRPO fine-tuning of the cold-started model in Stage 3 is composed of the Orsta47K (Ma et al., 2025) and virl39K (Wang et al., 2025a) datasets. In Stage 2 of cold start training, we used 9K self-distilled data. This composition is designed to enhance the model’s general and mathematical reasoning capabilities. We conduct evaluations on multiple benchmark datasets, including MEGA-Bench (Chen et al., 2025b), MMMU (Yue et al., 2024), MathVista (Lu et al., 2024b), MATH-Vision (Wang et al., 2024a), and MathVerse (Zhang et al., 2024).
@@
    The remaining results are from the MEGA-Bench Leaderboard and Ma et al. (2025).
@@
Baseline: Our comparative analysis is grounded on two primary categories of models. The first category comprises open-source general VLMs, including QwenVL-2-7B (Wang et al., 2024b), QwenVL-2.5-7B (Bai et al., 2025), InternVL2-8B (Chen et al., 2024), InternVL2.5-8B (Chen et al., 2024), Kimi-VL-A3B (Team et al., 2025), and DeepSeek-VL-7B (Lu et al., 2024a). The second category focuses on models specifically engineered for advanced reasoning tasks. This group includes Kimi-VL-A3B-Thinking (Team et al., 2025), R1-Onevision (Yang et al., 2025), VLAA-Thinking (Chen et al., 2025a), MM-Eureka-7B (Meng et al., 2025), VL-Rethinker-7B (Wang et al., 2025a), and Orsta-7B (Ma et al., 2025).
Implementation Details: We utilize the open-source Multimodal Large Language Model, Qwen2.5-VL-7B (Bai et al., 2025), as our base model. For the GRPO training in Stage 1 and Stage 3, we employ the MM-EUREKA  [^5] https://github.com/ModalMinds/MM-EUREKA framework. The rollout and training batch sizes are both set to $128$, with $8$ rollouts generated per sample. The learning rate is configured to $1\times 10^{-6}$. For the DPO training in Stage 2, as well as for the comparative SFT experiments, we leverage the LlamaFactory  [^6] https://github.com/hiyouga/LLaMA-Factory framework. In this configuration, the training batch size is set to $64$, the learning rate is maintained at $1\times 10^{-6}$, and the hyperparameter $\lambda$ for the hybrid loss function is set to $1$. The prompt used during training is shown in Appendix A.2.
@@ 4.3 Ablation on Self-Distillation and Decoupled Data Strategy
Self-distillation proves more effective than external teacher models. First, we evaluate the effectiveness of the self-distillation mechanism by substituting it with preference data generated from powerful external teacher models, specifically QwenVL-2.5-32B and QwenVL-2.5-72B. The results shown in Tabel 3 clearly indicate that our self-distillation approach outperforms both teacher-based alternatives. We also observe that performance degradation is more pronounced when using the QwenVL-2.5-32B model, whose output distribution diverges more significantly from our base model. This finding suggests that preference data closely aligned with the model’s intrinsic capability distribution is more effective for alignment than guidance from a more capable but dissimilar external model.

### v2

@@ 1 Introduction
Recently, inspired by the success of DeepSeek-R1 (guo2025deepseek), in effectively enhancing the reasoning capabilities of large models through reinforcement learning (RL) with verifiable reward (lambert2024tulu; guo2025deepseek), a growing body of work has begun to apply RL directly to vision language models (VLMs). This has led to a wave of exciting “MLLM-r1” research (meng2025mmeureka; shen2025vlm; peng2025lmm; zhou2025r1; zhang2025r1; wang2025sota; wang2025vl; zheng2025deepeyes; ma2025one), which leverage similar principles to advance multimodal reasoning.
@@
Previous research has indicated that prior to RL, employing a pre-training or warm-up phase (which is termed “cold start”), can significantly improve the readability, stability, and even the final performance of RL training (guo2025deepseek). Currently, the most commonly used cold start strategy is supervised fine-tuning (SFT) , where the model is first fine-tuned on a set of high-quality Chain-of-Thought reasoning data to provide a better initial policy for the subsequent RL phase (wei2025advancing; yang2025r1; huang2025vision; deng2025openvlthinker). This strategy enables the model to be trained on complex reasoning data during the cold start phase, thereby acquiring reasoning ability.
@@
The common understanding behind SFT-based cold start is that reasoning abilities, reasoning format and other learning objectives can be jointly learned during the cold start phase. However, such an SFT-based joint learning paradigm may largely affect the model’s generalization capability (wu2025generalization; chu2025sft), and consequently degrade subsequent RL (chen2025sft). This raises an important research issue of quantifying and improving the model’s generalization capability during cold start and working in concert with subsequent RL.
@@
To address the above limitations, we consider an alternative learning paradigm, which separates the learning process into hierarchical stages based on the idea that cold start phase focused more on shallow, surface-form learning to avoid prematurely getting stuck in in-distribution problem solving, while subsequent RL focuses on the deep-level learning of a solution to boost the overall performance (bengio2009curriculum). Thus, the intuition of our adopting decoupling learning for multimodal reasoning is that the selection of pre-training methods in cold start needs to better support the subsequent RL, both in terms of generalization and by having separate objectives to facilitate better final results.
@@
Another important issue is the generation of cold start data. Previously, the prohibitive cost of human annotation has motivated a growing body of research to explore the use of synthetic data. This often involves using a more capable large model as a “teacher” to distill data for a smaller “student” model. (zhang2025oasis; yao2024mulberry; xu2024llava; huang2025vision). However, when the capability gap between the teacher model and the student model is too large, it can lead to a decline in model performance (zhang2023towards). Alternatively, the DeepSeek-R1-Zero paradigm (guo2025deepseek) first directly applies reinforcement learning to the base model for obtaining R1-Zero and then generates cold start data by zero model itself. This paradigm has achieved very remarkable performance; yet it still has the limitation of reliance on the SFT cold start and the constraints between SFT and subsequent RL, thereby leaving room for further improvement.
@@
In this paper, to examine the suitable cold start training method, we propose the Generalization Factor (GF) coefficient in Section 2 to quantify the generalization capability of the model and conduct an empirical study to evaluate different training methods. We identify that Direct Preference Optimization (DPO) (rafailov2023direct) based on preference data is a cold start approach that enables the model to have better performance. On this basis, we present the Self-distilled Preference-based Cold-Start (SPECS) framework in Section 3. By decoupling the learning objectives during DPO to focus on output format, we create a pre-aligned model that serves as a superior starting point for the final RL fine-tuning. Our experiments show that this method leads to more stable, efficient training, and a higher performance ceiling compared to the advanced and strong baseline.
@@ 2.1 Evaluating Degree of Generalization
Setup. We define an evaluation function $\psi(f_{n},P)\in\mathbb{R}$ that measures the performance of a model $f$ on a data distribution $P$ and $n$ refers to the size of the training data samples. A higher value of $\psi$ indicates better performance.
Generalization Factor. To accurately evaluate the generalization ability of a model, we first need to test the model’s performance on in-distribution (ID) and out-of-distribution (OOD) tasks. Among them, ID tasks require the same as the task requirements during training, while OOD tasks require different from the task requirements during training.
  • ID Performance: $\Psi_{\mathrm{ID}}(n)$, is evaluated on a hold-out set from the same distribution $P_{train}$.
    $\Psi_{\mathrm{ID}}(n)=\psi(f_{n},P_{train})$
  • OOD Performance: $\Psi_{\mathrm{OOD}}(n)$, is the weighted average performance across a set of $m$ distinct OOD distributions, $Q=\{Q_{1},\ldots,Q_{m}\}$, with weights defined by a distribution $\alpha$.
    $\Psi_{\mathrm{OOD}}(n)=\mathbb{E}_{Q\sim\alpha}[\psi(f_{n},Q)]$
We establish a baseline model, $f_{0}$, which serves as a reference point. The performance gains over this baseline are calculated as:
$G_{\mathrm{ID}}(n)=\Psi_{\mathrm{ID}}(n)-\Psi_{\mathrm{ID}}(0)$
$G_{\mathrm{OOD}}(n)=\Psi_{\mathrm{OOD}}(n)-\Psi_{\mathrm{OOD}}(0)$
We define GF, $\Gamma(n)$ as the $F_{\beta}$-score of the model with respect to OOD performance gains and ID performance gains. The reason for adopting this metric is that the $F_{\beta}$-score is particularly suitable for average ratios. Its most prominent feature is that the result tends to lean toward the smaller number. This perfectly aligns with our needs: as long as either the ID or OOD performance is very poor, the final score will be very low. We can also control the size of $\beta$ to reflect the degree of importance we attach to OOD performance gains during the training process.
$\Gamma(n)=(1+\beta^{2})\frac{G_{\mathrm{ID}}(n)G_{\mathrm{OOD}}(n)}{\beta^{2}\cdot G_{\mathrm{ID}}(n)+G_{\mathrm{OOD}}(n)}$
@@ 2.2 Experimental findings
To preliminarily examine how preference data and supervised data affect model generalization, we construct a preference dataset $\mathcal{D}_{\mathrm{pref}}=\{(x_{i},y_{i}^{+},y_{i}^{-})\}_{i=1}^{N}$ , where $y_{i}^{+}$ is the chosen response and $y_{i}^{-}$ is the rejected response, and a supervised dataset $\mathcal{D}_{\mathrm{SFT}}=\{(x_{i},y_{i})\}_{i=1}^{N}$ with $y_{i}=y_{i}^{+}$ around reasoning tasks defined by a specific answer format. Under equal data budgets, we evaluate two settings: (i) an in-distribution setting in which the required reasoning format matches that used in training, and (ii) an out-of-distribution setting in which the required reasoning format differs  [^2] For more detailed description, see Appendix B. We compare DPO training, SFT training, and DPO training augmented with SFT loss (see Section 3.3). The resulting $\Psi_{\mathrm{ID}}(n)$, $\Psi_{\mathrm{OOD}}(n)$, and the $\Gamma(n)$ are reported in Figure 1.
@@ 3.2 Self-Distillation for Preference Data Generation
  • Chosen Response Filtration. For the chosen response ($y_{i}^{+}$), we use Gemini-2.5 flash (comanici2025gemini) as an evaluator. Assesses whether the reasoning path in the $\pi_{GRPO-zero}$ response aligns correctly with its final answer. Only responses in which the reasoning and the answer are consistent are retained, forming a high-quality pool of candidates. For more analysis of this content, please refer to Appendix E.
@@ 3.3 DPO-based Pre-Alignment for Cold-Start
Methodology: To achieve this pre-alignment, we employ DPO (rafailov2023direct), a powerful technique that directly optimizes the language model on preference data without the need for an explicit reward model. The standard DPO loss function is defined as:
@@
To augment this process, we incorporate an SFT loss computed on the “chosen” samples, which serves as a form of regularization. It ensures that while the model learns the directional preference signal from DPO, it does not drift far from the core distribution of high-quality text embodied by the chosen responses (rao2025apt). The combined loss function is thus:
@@
where $\mathcal{L}_{SFT}$ is the conventional negative log-likelihood loss on the chosen responses, and $\lambda$ is a weighting coefficient to balance the two objectives. For a discussion for $\lambda$, see Appendix C.
@@ 3.4 Final GRPO Fine-tuning
For the final stage of fine-tuning, we employ the GRPO algorithm (shao2024deepseekmath). This process is guided by a composite reward function that combines format and accuracy components to evaluate the model’s output, $o$, for a given question, $q$.
@@ 4.1 Experiment Settings
Dataset and Benchmark: The data utilized for training $\pi_{GRPO-zero}$ model in Stage 1 and for the final GRPO fine-tuning of the cold-started model in Stage 3 is composed of the Orsta47K (ma2025one) and virl39K (wang2025vl) datasets. In Stage 2 of cold start training, we used 9K self-distilled data. This composition is designed to enhance the model’s general and mathematical reasoning capabilities. We conduct evaluations on multiple benchmark datasets, including MEGA-Bench (chen2025mega-bench), MMMU (yue2024mmmu), MathVista (lu2024mathvista), MATH-Vision (wang2024measuring), and MathVerse (zhang2024mathverse).
@@
    The remaining results are from the MEGA-Bench Leaderboard and ma2025one.
@@
Baseline: Our comparative analysis is grounded on two primary categories of models. The first category comprises open-source general VLMs, including QwenVL-2-7B (Qwen2VL), QwenVL-2.5-7B (Qwen2.5-VL), InternVL2-8B (chen2024internvl), InternVL2.5-8B (chen2024internvl), Kimi-VL-A3B (kimiteam2025kimivltechnicalreport), and DeepSeek-VL-7B (lu2024deepseekvl). The second category focuses on models specifically engineered for advanced reasoning tasks. This group includes Kimi-VL-A3B-Thinking (kimiteam2025kimivltechnicalreport), R1-Onevision (yang2025r1), VLAA-Thinking (chen2025sft), MM-Eureka-7B (meng2025mmeureka), VL-Rethinker-7B (wang2025vl), and Orsta-7B (ma2025one).
Implementation Details: We utilize the open-source Multimodal Large Language Model, Qwen2.5-VL-7B (Qwen2.5-VL), as our base model. For the GRPO training in Stage 1 and Stage 3, we employ the MM-EUREKA  [^5] https://github.com/ModalMinds/MM-EUREKA framework. The rollout and training batch sizes are both set to $128$, with $8$ rollouts generated per sample. The learning rate is configured to $1\times 10^{-6}$. For the DPO training in Stage 2, as well as for the comparative SFT experiments, we leverage the LlamaFactory  [^6] https://github.com/hiyouga/LLaMA-Factory framework. In this configuration, the training batch size is set to $64$, the learning rate is maintained at $1\times 10^{-6}$, and the hyperparameter $\lambda$ for the hybrid loss function is set to $1$. The prompt used during training is shown in Appendix B. Some more detailed settings can be found in Appendix F.
@@ 4.3 Ablation on Self-Distillation and Decoupled Data Strategy
Self-distillation proves more effective than external teacher models. First, we evaluate the effectiveness of the self-distillation mechanism by substituting it with preference data generated from powerful external teacher models, specifically QwenVL-2.5-32B and QwenVL-2.5-72B. The results shown in Tabel 4 clearly indicate that our self-distillation approach outperforms both teacher-based alternatives. We also observe that performance degradation is more pronounced when using the QwenVL-2.5-32B model, whose output distribution diverges more significantly from our base model. This finding suggests that preference data closely aligned with the model’s intrinsic capability distribution is more effective for alignment than guidance from a more capable but dissimilar external model.
Distilling from GRPO-zero instead of the base model. We directly perform RL on the base model to obtain GRPO-zero, and then distill the chosen response through the GRPO-zero model. This choice of scheme is based on considerations of data utilization and training data quality. As shown in Table 3 below, we have conducted statistics on some indicators of the responses of the original model and the GRPO-zero model to the training data questions. Obviously, the GRPO-zero model has higher data utilization in the collection of chosen responses due to its higher format accuracy and answer accuracy. At the same time, we counted the number of reasoning words (including transition words, causal words, sequential words, etc.) per 1000 characters for both models. We also used the same data collection method to collect responses with correct formatting and correct answers distilled from the base model as chosen responses. The experimental results shown in Table 4.
Table 3: Statistical indicators of response in training data for the base model and GRPO-Zero model
Model Format Acc. (%) Answer Acc. (%) Reasoning Words / 1k Chars
Qwen2.5-7B-Instruct 41.62 30.42 4.26
Ours-GRPO-zero 96.74 52.82 4.99
@@
4.4 Analysis of the Impact of DPO-based Cold Start and SFT-based Cold Start
@@
Performance and Training Efficiency. To evaluate performance and training efficiency, we tracked MEGA-Bench scores throughout the GRPO training process. As illustrated in Figure 4, the DPO-based GRPO model begins with a substantially higher initial score, demonstrating the immediate benefit of preference-based pre-alignment. Furthermore, it maintains a clear advantage throughout training, converging more rapidly and ultimately achieving a higher performance ceiling than its SFT-based GRPO counterpart. The final performance comparison of other benchmarks is shown in Table 5 below. For more analysis on this content, please refer to Appendix D.
Table 5: Performance Comparison of SFT-based, and DPO-based Models on Various Benchmarks
Model Megabench MMMU MathVista MathVision MathVerse AVG
Qwen2.5-7B-Instruct 35.07 54.20 63.70 25.40 38.20 43.31
SFT-based GRPO 37.52 54.44 74.10 28.61 43.60 47.65
DPO-based GRPO 39.17 56.78 75.90 29.50 48.73 50.02
Training Stability. Beyond performance metrics, we analyzed training stability by comparing the policy loss curves, presented in Figure 4. The curve for DPO-based GRPO is visibly smoother and more stable, indicating a more consistent and reliable optimization trajectory. In contrast, the SFT-based GRPO policy exhibits greater volatility, suggesting that the RL algorithm make more drastic and potentially erratic updates. In terms of format rewards, RL based on SFT cold start is also weaker than RL based on DPO cold start in terms of the stability of format rewards. Regarding the impact of different cold-start training methods on the stability of model training, we believe this is related to the training objectives of the cold-start phase and the RL phase. The SFT training objective is to maximize log likelihood, which is a form of imitation learning, while the loss function of DPO can be seen as directly optimizing an implicit reward model consistent with preference data, which is more aligned with the subsequent reward-driven GRPO optimization objective. Therefore, using a DPO-based model as a starting point also brings more stable training for subsequent RL.